Mary mixes white paint and blue paint in the ratio 2:3.
She makes a total of 20 litres of paint.
How much more blue paint does she need to add to the mixture so that the ratio of white paint to blue paint becomes 1:7?

Answer by ikleyn(52799)

.
20 liters of paint consist of 2+3 = 5 equal parts; so each part is 4 liters; hence, white paint is 8 liters  (and blue paint is 12 liters).


To get the final ratio white to blue as 1:7, Mary must have 7*8 = 56 liters of the blue paint - so she need to add 56-12 = 44 liters.


Answer.  44 liters of the blue paint to add.